OpenStudy (anonymous):

A metal worker has several 1-kilogram bars of a metal alloy that contain 25% copper and several 1-kilogram bars that contain 61% copper. How many bars of each type of alloy should be melted and combined to create 54 kilograms of a 49% copper alloy? Set up and solve a system of equations to find the answer.

OpenStudy (aum):

Let number of 25% copper bars be x and number of 61% copper bars be y. They are combined to form 54 kilograms of a 49% copper alloy. x + y = 54 0.25x + 0.61y = 54 * 0.49 (copper content) Solve the two equations.

OpenStudy (anonymous):

x=18 and y=36?

OpenStudy (anonymous):

thanks! @aum